OpenStudy (anonymous):

Same problem with graphing, once I have found my solution set. x-3y=-3 x+3y=9 I get (3, 2) as my solution How do I graph this?

OpenStudy (anonymous):

you just put what ever equation in slope intercept form graph

OpenStudy (anonymous):

example x-3y=-3 -3y = -x -3 -1/3 [ -3 y = -x -3 = y = (1 x / 3 ) + 1 slope = 1/3 y intercept = 1

OpenStudy (anonymous):

2 points determine a line. given the y-intercept , for the first equation which is 1, we can create an ordered pairs (0,1) where x at 0 that,s for the first equation. given the solution (3,2) and another point (0,1), graph it. . for equation 2, same process with equation 1.

Directrix (directrix):

@ LynDor --> are you wanting to graph the system of equations x-3y=-3, x+3y=9? Or, do you want to graph the line containing the point of intersection (common solution) of the two given equations?

Directrix (directrix):

Clink this link for the graphs of the system of equations: http://www.wolframalpha.com/input/?i=%28x-3y%3D-3%29+and+%28x%2B3y%3D9%29 Infinitly many lines can be drawn that contain the point (3,2).

OpenStudy (anonymous):